OpenStudy (anonymous):

What is the solution of the equation: 5/8x - 3= 8

OpenStudy (anonymous):

You want just the solution or the process?

OpenStudy (anonymous):

x=88 over 5

OpenStudy (anonymous):

the process would be nice too

OpenStudy (austinl):

\(\large{\frac{5}{8}x-3=8}\) \(\large{\frac{5}{8}x=11}\) \(\large{x=11\times\frac{8}{5}}\)

OpenStudy (austinl):

Did that make sense?

OpenStudy (anonymous):

yes thank you so much

OpenStudy (austinl):

No problemo! :)